# arXiv:2603.00822v1[cs.SE] 28 Feb 2026

### 28 Feb 2026 v1

@@ 3.2. Constraint Extraction
The extraction phase transforms unstructured Markdown into a set of refined, standalone constraints. This is challenging because Agent Instructions are written for human readers who naturally understand context and scope from document structure.
The Challenge of Lexical Scope.
Simple approaches to Agent Instruction segmentation fail to preserve lexical scope. Consider the following excerpt:
@@
The instruction “Use pytest” appears as a standalone bullet, but its meaning depends on its hierarchical context and applies to Backend Testing, not to the entire project or to Database operations. If we extract this constraint without its context, we might incorrectly enforce pytest usage for frontend code or database migrations.
Markdown Abstract Syntax Tree.
To preserve this hierarchical context, we parse each Agent Instruction file into a tree structure. Internal nodes represent headers at various levels (H1 through H6), leaf nodes represent content blocks (paragraphs, list items, code blocks), and edges represent containment relationships. This tree structure mirrors how humans mentally organize the document when reading it.
Path-Aware Slicing.
For each leaf node representing a potential constraint, we compute its header path, the ordered sequence of ancestor headers from the document root to the leaf. In the example above, “Use pytest” has the header path [Backend, Testing]. We then use an LLM to rewrite the raw instruction into a standalone, unambiguous statement that incorporates this context, producing a refined constraint like “For backend testing, use the pytest framework for all unit and integration tests.” This refinement step disambiguates scope and normalizes diverse phrasings into consistent constraint statements that can be enforced without additional context.
Stable Identifiers for Incrementality.
To support incremental updates, we assign each constraint a stable identifier by hashing its header path and content. When a developer modifies one section of an Agent Instruction file, only constraints from that section need to be re-extracted and re-synthesized. Constraints from unmodified sections retain their identifiers and their previously generated checks remain valid. This is important for large instruction files where full re-extraction would be expensive.
@@ 3.3. Constraint to Check
This phase transforms refined constraints into executable Python code. Rather than using a single monolithic code generator, we route each constraint to a specialized generator based on its enforcement domain. This domain-specific approach improves code quality because each generator is optimized for its particular verification pattern.
The Intent Router.
The Router classifies each constraint into one of four enforcement domains. For the Process Domain, the router extracts target binary names (e.g., npm, yarn), the enforcement level, and the applicable scope. For the Source Domain, it identifies target programming languages, the pattern type to detect, and the severity level. For the Architectural Deterministic Domain, it determines the constraint type (dependency direction, cycle detection, or boundary enforcement) and selects an appropriate graph algorithm. For the Architectural Semantic Domain, it preserves the natural language description of the design principle for later LLM evaluation.
The router’s classification is performed by the LLM, prompted with examples of each domain to improve accuracy.
Specialized Code Generators.
Each domain has a specialized Expert Generator that produces Python code tailored to that domain’s verification needs. The Process Expert generates code that inspects command-line arguments and environment variables, returning a verdict that determines whether the command should proceed. The Source Expert generates code using Tree-sitter bindings that constructs AST queries and traverses parse trees, returning violations with precise file and line information. The Architectural Expert generates code using NetworkX that applies graph algorithms to dependency structures, detecting cycles, layer violations, or boundary crossings.
Tree-sitter grammars are bundled for Python, TypeScript, JavaScript, Go, and Rust. This coverage addresses the most common languages in repositories with Agent Instructions, though additional languages can be added by installing the corresponding Tree-sitter packages.
Check Storage.
Synthesized checks are stored in a JSON file that can be reviewed and manually edited by developers before deployment, providing a layer of human oversight over the generated checks. The listing below shows an example entry:
"constraints": {
"a7f3b2...": {
"domain": "PROCESS",
"original_text": "NEVER use npm run compile",
"refined_text": "Block npm run compile commands",
"source_file": "copilot-instructions.md",
"header_path": ["Validating TypeScript changes"],
"enforcement_level": "block",
"check_code": "def check(cmd, args, env): ..."
@@ 3.4. Runtime Check Enforcement
The enforcement phase executes synthesized checks through four specialized modules, each designed for a particular type of constraint.
Process Interceptor.
Enforcing process constraints requires intercepting shell commands before execution. We use \$PATH manipulation for its simplicity and portability. ContextCov creates a shim directory containing wrapper scripts for monitored binaries and prepends this directory to \$PATH. Each monitored binary (e.g., npm, yarn, pip) has a corresponding wrapper script:
@@
exec python -m contextcov.shim npm "\$@"
@@
Universal Static Linter.
ContextCov uses Tree-sitter as a universal parser for source code analysis. Tree-sitter provides incremental parsing, graceful error recovery, and a uniform API across languages, making it ideal for analyzing code in diverse repositories. The linter scans source files matching configured patterns, parses each file with the appropriate grammar based on file extension, and runs all applicable checks. Each check receives the parse tree, the raw source bytes, and the file path, allowing it to construct queries, traverse the AST, and report violations with precise locations. The following example shows a generated check that detects Promise.then() usage:
@@
"""CheckforPromise.then()usage."""
query = LANGUAGE.query("""
␣␣␣␣␣␣␣␣(call_expression
␣␣␣␣␣␣␣␣␣␣function:(member_expression
␣␣␣␣␣␣␣␣␣␣␣␣property:(property_identifier)@method)
␣␣␣␣␣␣␣␣␣␣(#match?@method"^then\$"))""")
@@
"Useasync/awaitinsteadof.then()"))
@@
Architectural Validator.
Deterministic architectural constraints are enforced by constructing and querying a dependency graph. The graph is built by parsing import statements using language-appropriate techniques. Python’s ast module extracts import and from ... import statements, Tree-sitter queries extract import declarations and require() calls from JavaScript and TypeScript, and regex-based extraction handles Go import blocks. The resulting graph is stored as a NetworkX DiGraph where nodes represent modules and edges represent dependencies.
Architectural checks operate on this graph to detect violations such as layer crossings, cycles, or improper module placement. The following example enforces that files under src/vs/workbench/ must reside in one of the allowed subdirectories:
@@
"""Enforceworkbench/subdirectorystructure."""
allowed = {’browser’, ’services’, ’contrib’, ’api’}
@@
message=f"Mustbein{allowed}"))
@@
Semantic Architectural Validator.
Some architectural constraints cannot be verified through static analysis alone because they require understanding the intent behind code. For example, the constraint “Do not use another component’s storage keys, define a proper API instead” cannot be checked by examining import graphs or file paths. It requires understanding what constitutes a “storage key,” which code belongs to which “component,” and whether an access pattern constitutes a violation.
For these semantic constraints, we use an LLM-as-judge approach (Crupi et al., 2025). When evaluating code changes, the LLM receives the constraint text, the relevant code snippet, and surrounding context such as file path and neighboring code. The LLM then judges whether the code respects the design principle, providing a verdict and explanation. Semantic checks produce WARNING verdicts rather than hard blocks, flagging potential violations for human review while avoiding false positives from overly strict interpretation.
@@ 3.5. Handling Ambiguity
We adopt a fail-closed philosophy where ambiguous constraints are interpreted strictly. “Use pnpm” triggers blocking of npm, yarn, and bun globally, even though a human reader might interpret this more narrowly. This approach may cause false positives, but we argue this is the correct trade-off. False positives can be overridden by developers who understand the context, either by refining the Agent Instructions to be more precise or by marking specific checks as exceptions. False negatives, on the other hand, allow violations to slip through undetected, accumulating as technical debt. This philosophy aligns with security principles where, when uncertain, the system denies the action and requires explicit authorization.
@@ 4. Evaluation
We evaluate ContextCov on a large corpus of open-source repositories containing Agent Instruction files. Our evaluation addresses two key questions. First, can ContextCov reliably extract constraints from diverse Agent Instructions and synthesize them into valid, executable checks (RQ1)? Second, when applied to real codebases, does ContextCov detect meaningful violations that would otherwise go unnoticed (RQ2)?
@@ 4.1. Experimental Setup
We use a subset of the dataset from Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025), which contains 1,925 GitHub repositories with Agent Instruction files. From this dataset, we selected the top 723 repositories ranked by star count, with the highest-starred repository having 182k stars. Table 1 summarizes the dataset characteristics. All experiments used gpt-5.2-chat for LLM-based operations and ran on an Intel Xeon Platinum 8160M CPU with 754GB RAM running Rocky Linux 9.7. We used Tree-sitter (35) for AST parsing with grammars for Python, TypeScript, JavaScript, Go, and Rust.
Table 1. Dataset statistics.
Repositories Median Stars Median Lines Median Words
723 2,010 98 603
4.2. RQ1. Extraction and Synthesis Quality
We evaluate how effectively ContextCov extracts constraints from Agent Instructions and synthesizes them into valid executable checks. Figure 3 visualizes the extraction pipeline flow across all 723 repositories.
Figure 3. Extraction pipeline flow across 723 repositories. The pipeline processed 51,490 AST leaf nodes, extracted 48,921 non-empty segments, and synthesized 46,316 checks across four enforcement domains. Source accounts for 28%, Process for 26%, Architectural Deterministic for 20%, and Architectural Semantic for 26%.
A key concern with LLM-based code generation is whether the synthesized checks are syntactically valid and executable. We validated all generated checks by parsing them and checking for syntax errors. Of the 34,374 checks that require syntax validation (Source, Process, and deterministic Architectural checks), 34,373 passed validation (99.997%). Only 1 check failed due to a syntax error. Semantic Architectural checks are excluded from this count as they use natural language prompts rather than executable code. This high validity rate demonstrates that ContextCov’s domain-specialized code generators produce reliable, executable checks with minimal need for manual intervention. On average, each repository produces 64 executable checks after filtering and synthesis.
4.3. RQ2. Detection in the Wild
We evaluate ContextCov’s ability to detect constraint violations in real-world repositories. In practice, ContextCov would run on code diffs as an agent makes changes, flagging violations before they are committed. For evaluation purposes, we run ContextCov against entire repositories to assess its detection capability. This approach is necessary because these repositories do not contain metadata identifying which code was contributed by agents. It is also meaningful because agents learn from existing code patterns. Any violation found in the existing codebase represents a pattern that an agent could mimic or propagate, making repository-wide analysis a proxy for the violations agents are likely to introduce.
Table 2 reports violation statistics across all 723 repositories. Of the 46,316 synthesized checks, 10,927 (24%) detected at least one violation.
Table 2. Violation detection statistics across all repositories.
Domain Total Checks Checks w/ Violations
Source 13,183 4,232 (32%)
Process 12,080 3,768 (31%)
Arch (det) 9,111 2,605 (29%)
Arch (sem) 11,942 322 (3%)
Total 46,316 10,927 (24%)
Figure 4 shows how violations are distributed across repositories by domain. Figure 5 shows that violations follow a long-tail distribution, highlighting areas where agents are most likely to introduce drift if not guided by enforcement.
Figure 4. Distribution of violations across repositories by domain. The distribution exhibits high variance, with some repositories having thousands of violations while others have none, reflecting diversity in Agent Instruction specificity and codebase compliance.
Figure 5. Violation concentration across checks. A small number of constraints account for the majority of violations, suggesting that certain constraint types are systematically violated across codebases.
Overall, ContextCov detects over 500,000 violations across 723 repositories, with 81% of repositories containing at least one violation. This demonstrates that Context Drift between documented Agent Instructions and actual code is widespread, validating the need for automated enforcement to bridge the Reality Gap.

### 04 May 2026 v2

@@ 3.2. Constraint Extraction
The extraction phase transforms unstructured Markdown into a set of refined, standalone constraints. This is challenging because Agent Instructions are written for human readers who naturally understand context and scope from document structure. Simple approaches to segmentation fail to preserve lexical scope. Consider an excerpt where “Use pytest” appears under the headers “Backend” and “Testing”. The instruction’s meaning depends on its hierarchical context as it applies to Backend Testing, not to the entire project or to Database operations. Extracting this constraint without its context could incorrectly enforce pytest usage for frontend code or database migrations.
@@
To preserve hierarchical context, we parse each Agent Instruction file into a tree structure where internal nodes represent headers at various levels (H1 through H6), leaf nodes represent content blocks (paragraphs, list items, code blocks), and edges represent containment relationships. This tree mirrors how humans mentally organize the document when reading it. For each leaf node representing a potential constraint, we compute its header path—the ordered sequence of ancestor headers from the document root to the leaf. In the example above, “Use pytest” has the header path [Backend, Testing]. We then use an LLM to rewrite the raw instruction into a standalone, unambiguous statement that incorporates this context, producing a refined constraint like “For backend testing, use the pytest framework for all unit and integration tests.” This refinement step disambiguates scope and normalizes diverse phrasings into consistent constraint statements that can be enforced without additional context.
To support incremental updates, we assign each constraint a stable identifier by hashing its header path and content. When a developer modifies one section of an Agent Instruction file, only constraints from that section need to be re-extracted and re-synthesized. Constraints from unmodified sections retain their identifiers and their previously generated checks remain valid, which is important for large instruction files where full re-extraction would be expensive.
@@ 3.3. Constraint to Check
This phase transforms refined constraints into executable Python code. Rather than using a single monolithic code generator, we route each constraint to a specialized generator based on its enforcement domain. This domain-specific approach improves code quality because each generator is optimized for its particular verification pattern. The Router classifies each constraint into one of four enforcement domains: the Process Domain extracts target binary names (e.g., npm, yarn), the enforcement level, and the applicable scope; the Source Domain identifies target programming languages, the pattern type to detect, and the severity level; the Architectural Deterministic Domain determines the constraint type (dependency direction, cycle detection, or boundary enforcement) and selects an appropriate graph algorithm; and the Architectural Semantic Domain preserves the natural language description of the design principle for later LLM evaluation. The router’s classification is performed by the LLM, prompted with examples of each domain to improve accuracy.
Each domain has a specialized Expert Generator that produces Python code tailored to that domain’s verification needs. The Process Expert generates code that inspects command-line arguments and environment variables, returning a verdict that determines whether the command should proceed. The Source Expert generates code using Tree-sitter bindings that constructs AST queries and traverses parse trees, returning violations with precise file and line information. The Architectural Expert generates code using NetworkX that applies graph algorithms to dependency structures, detecting cycles, layer violations, or boundary crossings. In our implementation, domain-specific outputs are generated through Claude Code with Opus 4.5 rather than by a single direct model call. The generation prompt provides the refined constraint, repository structure, and relevant files, and instructs an iterative self-test loop that samples representative outcomes and revises logic when evidence suggests false positives. Our pipeline also enforces programmatic validation for generated executable checks through syntax checks and dry-run execution before accepting the check, while semantic architectural generation produces reviewer rubrics rather than executable code and uses output-shape validation instead. Tree-sitter grammars are bundled for Python, TypeScript, JavaScript, Go, and Rust, covering the most common languages in repositories with Agent Instructions.
Synthesized checks are stored in a JSON file that can be reviewed and manually edited by developers before deployment, providing a layer of human oversight over the generated checks. The listing below shows an example entry.
constraints”: {
a7f3b2…”: {
domain”: PROCESS”,
original_text”: NEVER use npm run compile”,
refined_text”: Block npm run compile commands”,
source_file”: copilot-instructions.md”,
header_path”: [”Validating TypeScript changes”],
enforcement_level”: block”,
check_code”: def check(cmd, args, env): …”
@@ 3.4. Runtime Check Enforcement
The enforcement phase executes synthesized checks through four specialized modules, each designed for a particular type of constraint. Enforcing process constraints requires intercepting shell commands before execution, which we accomplish through \$PATH manipulation for simplicity and portability. ContextCov creates a shim directory containing wrapper scripts for monitored binaries and prepends this directory to \$PATH. Each monitored binary (e.g., npm, yarn, pip) has a corresponding wrapper script:
@@
exec python -m contextcov.shim npm \$@
@@
For source code analysis, ContextCov uses Tree-sitter (Tree-sitter, [n. d.]), which provides incremental parsing, graceful error recovery, and a uniform API across languages. These properties make it suitable for analyzing code in diverse repositories. The linter scans source files matching configured patterns, parses each file with the appropriate grammar based on file extension, and runs all applicable checks. Each check receives the parse tree, the raw source bytes, and the file path, allowing it to construct queries, traverse the AST, and report violations with precise locations. The following example shows a generated check that detects Promise.then() usage:
@@
”””Check for Promise.then() usage.”””
query = LANGUAGE.query(”””
(call_expression
function: (member_expression
property: (property_identifier) @method)
(#match? @method ”^then\$”)) ”””)
@@
Use async/await instead of .then()”))
@@
Deterministic architectural constraints are enforced by constructing and querying a dependency graph. The graph is built by parsing import statements using language-appropriate techniques: Python’s ast module extracts import and from ... import statements, Tree-sitter queries extract import declarations and require() calls from JavaScript and TypeScript, and regex-based extraction handles Go import blocks. The resulting graph is stored as a NetworkX DiGraph where nodes represent modules and edges represent dependencies. Architectural checks operate on this graph to detect violations such as layer crossings, cycles, or improper module placement. The following example enforces that files under src/vs/workbench/ must reside in one of the allowed subdirectories:
@@
”””Enforce workbench/ subdirectory structure.”””
allowed = {’browser’, services’, contrib’, api’}
@@
message=fMust be in {allowed}”))
@@
Some architectural constraints cannot be verified through static analysis alone because they require understanding the intent behind code. For example, the constraint “Do not use another component’s storage keys, define a proper API instead” cannot be checked by examining import graphs or file paths—it requires understanding what constitutes a “storage key,” which code belongs to which “component,” and whether an access pattern constitutes a violation. For these semantic constraints, we use an LLM-as-judge approach (Crupi et al., 2025). When evaluating code changes, the LLM receives the constraint text, the relevant code snippet, and surrounding context such as file path and neighboring code. The LLM then judges whether the code respects the design principle, providing a verdict and explanation. Semantic checks produce WARNING verdicts rather than hard blocks, flagging potential violations for human review while avoiding false positives from overly strict interpretation.
@@ 3.5. Handling Ambiguity
We adopt a fail-closed philosophy where ambiguous constraints are interpreted strictly. “Use pnpm” triggers blocking of npm, yarn, and bun globally, even though a human reader might interpret this more narrowly. This approach may cause false positives, but we argue that this is the better trade-off for autonomous operation. False positives can be overridden by developers who understand the context, either by refining the Agent Instructions to be more precise or by marking specific checks as exceptions. False negatives allow violations to slip through undetected and accumulate as technical debt. This philosophy aligns with security principles where the system denies the action when uncertainty remains and requires explicit authorization (Saltzer and Schroeder, 1975).
@@ 4. Evaluation
We evaluate ContextCov on the SWE-bench Lite dataset to assess its extraction accuracy and its dynamic impact on autonomous agents. Our evaluation is driven by four research questions:
  • RQ1 (Extraction Accuracy): How accurately does ContextCov translate natural-language Agent Instructions into semantically correct executable constraints?
  • RQ2 (Enforcement Effectiveness): To what extent does active executable feedback reduce constraint violations compared to LLM reflection?
  • RQ3 (Functional Correctness): Does enforcing strict constraints negatively impact the agent’s ability to functionally resolve issues?
  • RQ4 (Overhead): What is the relative cost and latency of executable feedback compared to LLM reflection?
Aggregated statistics are derived from 300 SWE-bench Lite tasks executed across three agent modes.
@@ 4.1. Experimental Setup
We evaluate on the 12 repositories comprising the 300 tasks in SWE-bench Lite: astropy, django, matplotlib, seaborn, flask, requests, xarray, pylint, pytest, scikit-learn, sphinx, and sympy. Because historical SWE-bench issues predate modern agentic workflows, we systematically generated repository-grounded AGENTS.md files for all 12 repositories. To ensure construct validity, we used a few-shot LLM generator to extract explicit constraints strictly from existing developer documentation (e.g., CONTRIBUTING.md), linting configurations (pyproject.toml, setup.cfg, .pre-commit-config.yaml), and CI workflows (.github/workflows/, pytest.ini, tox.ini). We strictly prompted the generator against hallucination, requiring all rules to be supported by existing repository artifacts.
Given these instruction files, ContextCov generated domain-specific checks (source, process, architectural deterministic, and architectural semantic) and executed them across each repository to identify legacy violations and during the benchmark run to capture live violations.
We use OpenCode (Anomaly, [n. d.]), an open-source agentic coding framework, as our agent scaffold. The coding LLM is GPT-5.2 and the LLM Reflection critic uses Claude Opus 4.5. We run all 300 tasks from SWE-bench Lite across three agentic configurations:
  • Vanilla (baseline): the agent receives AGENTS.md in context but no active execution-time compliance feedback.
  • LLM Reflection (baseline): a critic LLM reviews candidate patches against AGENTS.md and returns natural-language feedback in a macro-loop.
  • ContextCov: the same loop structure, but feedback is produced by executable checks with deterministic violation traces.
Feedback-conditioned runs (LLM Reflection and ContextCov) share a maximum correction budget of five rounds.
4.2. RQ1: Extraction and Synthesis Quality
Table 1 (left) summarizes extraction and synthesis statistics across the 12 repositories. ContextCov synthesized 788 executable checks from instruction files across all four domains. Nearly all generated Python checks (780/781) parsed successfully, indicating robust syntactic generation.
To evaluate semantic correctness, we ran checks across repositories, identifying 5,435 potential legacy violations. We uniformly sampled 384 violations (95% confidence, $\pm$4.72% margin) for blind human annotation, achieving 82.81% extraction precision. An LLM-as-judge (Claude Opus 4.5) achieved 85.42% agreement with human annotators (Cohen’s $\kappa=0.61$), validating its use for subsequent dynamic evaluations.
Table 1. Repository-level statistics: extraction & check synthesis (RQ1, left) and harness resolution (RQ3, right). Van. = Vanilla, LR = LLM Reflection, $C$ = ContextCov.
Extraction & Synthesis (RQ1) Harness Resolution (RQ3)
Repository Lines Seg. Src Proc A-D A-S Chk $n$ Van. % LR $C$ $\Delta$
astropy 151 54 25 17 15 12 69 6 3 50 3 2 1-1
django 110 61 35 22 5 11 73 114 68 60 70 77 +9+9
matplotlib 157 71 30 24 20 9 83 23 8 35 10 9 +1+1
seaborn 106 44 15 23 13 0 51 4 2 50 2 2 0
flask 102 52 25 38 15 1 79 3 1 33 1 1 0
requests 100 54 16 30 15 6 67 6 6 100 2 6 0
xarray 44 17 4 12 1 2 19 5 1 20 2 2 +1+1
pylint 255 130 13 76 27 19 135 6 3 50 2 0 3-3
pytest 156 55 29 19 13 6 67 17 9 53 9 10 +1+1
scikit-learn 23 12 2 2 0 5 9 23 16 70 13 15 1-1
sphinx 104 43 13 17 14 9 53 16 6 38 5 7 +1+1
sympy 190 72 34 16 20 13 83 77 36 47 38 41 +5+5
Total 1,498 665 241 296 158 93 788 300 159 53 157 172 +13+13
4.3. RQ2: Effectiveness of Active Constraint Enforcement
RQ2 asks: To what extent does active, executable feedback reduce constraint violations and improve correction efficiency compared to LLM reflection?
We define a final clean patch as one containing zero actionable violations upon completion. For ContextCov, violations are intercepted dynamically in-loop; for baselines, we apply the executable oracle post-hoc. ContextCov achieves 88.3% compliance compared to 67.0% (vanilla) and 50.3% (LLM Reflection). Notably, LLM Reflection performs worse than the vanilla agent, suggesting that LLM-based critique introduces more drift than it corrects. A post-hoc taxonomy analysis revealed that all of ContextCov’s residual violations were limited to source-level checks and the process and architectural violations were fully eliminated by deterministic feedback.
Table 2. RQ2: Enforcement efficiency—oracle compliance, feedback rounds, and post-hoc LLM critic analysis (300 instances).
Oracle Compliance & Feedback Rounds Post-hoc LLM Critic
[…]
ContextCov 265 88.3% 0.14 1.67 1.31 29 (10%) ContextCov finals 30.1% (88/292) 292 1.15 336
Vanilla 201 67.0% 0.62 Vanilla finals 29.2% (87/298) 298 1.15 344
LLM Reflection 151 50.3% 1.01 2.85 2.30 85 (28%) LLM Refl. finals 50.3% (151/300) 300 1.01 304
ContextCov also requires fewer feedback rounds to converge (Table 2). LLM Reflection frequently exhausts the five-round cap, exhibiting “hallucination looping” where the LLM repeatedly fails to address the same issue. Deterministic error traces allow the agent to pinpoint violations faster. On instances where ContextCov and LLM Reflection disagreed on final outcomes, ContextCov won nearly twice as often (Figure 4), confirming that strict enforcement yields better patch trajectories.
Figure 3. RQ2: Harness resolution rate versus number of feedback rounds.
Figure 4. RQ2: Head-to-head wins when one method resolves an instance another does not.
We also applied the LLM Reflection critic post-hoc to all final patches (Table 2, right). The critic rates Vanilla and ContextCov finals similarly, suggesting its complaints are orthogonal to the constraints ContextCov enforces. Interestingly, LLM Reflection’s own finals achieve the highest critic-clean rate, yet the lowest harness resolution (RQ3). This disconnect suggests LLM critics may fixate on stylistic preferences unrelated to functional correctness.
4.4. RQ3: Impact on Functional Correctness
RQ3 asks: Does enforcing strict contextual constraints negatively impact the autonomous agent’s ability to successfully resolve functional software issues?
We applied the official SWE-bench Lite evaluation harness to final patches from each method. ContextCov resolves 57.3% of instances compared to 53.0% (vanilla) and 52.3% (LLM Reflection). McNemar’s tests on paired outcomes reject the null hypothesis of equal failure rates ($p<0.05$ for both comparisons), demonstrating that strict constraint enforcement improves rather than hinders functional correctness.
Table 3. RQ3: Functional correctness—harness resolution, pairwise wins, and McNemar tests (n=300n{=}300).
Method Resolved Rate vs Van. vs other FB Unique
ContextCov 172 57.3% +22 +32 9
LLM Reflection 157 52.3% +23 +17 10
Vanilla 159 53.0% +25 2
McNemar paired tests (discordant cells: A only / B only)
Vanilla vs. ContextCov: 9 / 22 $p=0.031$
Vanilla vs. LLM Reflection: 25 / 23 $p=0.885$
LLM Reflection vs. ContextCov: 17 / 32 $p=0.046$
Table 3 also reports pairwise wins and unique successes. Statistical significance arises because ContextCov’s wins outnumber its losses, not from isolated outliers. Per-repository analysis (Table 1, right) shows gains concentrated in larger subsets (django, sympy).
To understand failure modes, we cross-tabulated oracle cleanliness against harness outcomes (Table 4). For ContextCov, failures are largely independent of constraint violations and when it fails, the agent simply failed to deduce the bug’s logic, not because constraints trapped it. For baselines, oracle-dirty patches correlate with harness failures, suggesting uncorrected constraint violations contribute to functional failure.
Table 4. RQ3: Oracle cleanliness $\times$ harness resolution (n=300n{=}300 per method).
Method Oracle Pass Fail Total
ContextCov Clean 156 109 265
Dirty 16 19 35
LLM Refl. Clean 85 66 151
Dirty 72 77 149
Vanilla Clean 110 91 201
Dirty 49 50 99
Finally, we analyzed the overlap in success across the three methods. Of 300 instances, 127 passed under all three configurations, while 109 failed across the board. Notably, at least one method passed on 191 instances, yielding an implied union success rate of 63.7%, indicating that while ContextCov is the strongest individual method, different feedback mechanisms occasionally shepherd the agent down disparate, successful architectural paths.
4.5. RQ4: Overhead and Cost-Efficiency
[…]
Table 5. RQ4: Feedback cost proxies (300 instances).
Metric Method Mean Sum
LLM input chars LLM Refl. 19,455 5.84M
ContextCov 6,292 1.89M
Feedback time (s) LLM Refl. 133.7 40,110
ContextCov 27.7 8,310
Mean Median
End-to-end (s) LLM Refl. 461.4 379.2
ContextCov 422.4 240.0
Vanilla 327.9 184.7
Figure 5. RQ4: Mean logged LLM input characters (left) and mean feedback-generation wall time (right) for LLM Reflection versus ContextCov.
[…]